peanuts cost $2.00 per pound and cashews cost $6.00 per pound. how many pounds of each nut should be combined to get a five pound mixture at $4.40 per pound?
Answer by stanbon(75887)

peanuts cost $2.00 per pound and cashews cost $6.00 per pound. how many pounds of each nut should be combined to get a five pound mixture at $4.40 per pound?
----
Equations:
Quantity: p + c = 5 lbs
Value::: 2p +6c = 4.4*5 dollars
-------------------
Modify:
p + c = 5
p + 3c = 2.2*5
-----------------------
Subtract and solve for "c":
2c = 6
c = 3 lbs (amt. of cashews needed)
----
Solve for "p":
p + c = 5
p = 2 lbs (amt. peanuts needed)
=======================================================
Cheers,
Stan H.